OpenStudy (anonymous):

Someone tell me what I did wrong? I got this question wrong on a test and I am wondering how I should have done it differently. Simplify the expression. (3 + 2i)(4 – i) 12-3i+8i-2i^2 12+5i-2i^2 -2i^2+5i+12

OpenStudy (anonymous):

you can still simplify further: \[i^2 = -1\]

OpenStudy (anonymous):

Ohhh, I see, thanks.

OpenStudy (anonymous):

anytime you have the imaginary constant i raised to an even exponent it takes on a numerical value which is real

OpenStudy (anonymous):

\[i^1=i\] \[i^2=-1\] \[i^3 = -i\] \[i^4= 1\] i, -1, -i, 1 repeated for higher exponents